### v1

# Beyond the Use-and-then-Forget (UatF) Bound: Fixed Point Algorithms for Statistical Max-Min Power Control

Abstract
We introduce mathematical tools and fixed point algorithms for optimal statistical max-min power control in cellular and cell-less massive MIMO systems. Unlike previous studies that rely on the use-and-then-forget (UatF) lower bound on Shannon achievable (ergodic) rates, our proposed framework can deal with alternative bounds that explicitly consider perfect or imperfect channel state information (CSI) at the decoder. In doing so, we address limitations of UatF-based algorithms, which inherit the shortcomings of the UatF bound. For example, the UatF bound can be overly conservative: in extreme cases, under fully statistical (nonadaptive) beamforming in zero-mean channels, the UatF bound produces trivial (zero) rate bounds. It also lacks scale invariance: merely scaling the beamformers can change the bound drastically, especially when simple beamforming strategies are employed. In contrast, our framework is compatible with information-theoretic bounds that do not suffer from the above drawbacks. We illustrate the framework by solving a max-min power control problem considering a standard bound that exploits instantaneous CSI at the decoder.
I Introduction
The use-and-then-forget (UatF) bound has been fundamental in the development of beamforming and power control algorithms in massive MIMO and cell-less networks [1, 2, 3, 4, 5, 6, 7] because it often leads to tractable and scalable optimization problems. Furthermore, in some scenarios of practical relevance – e.g., when optimal beamformers are employed – the UatF bound often provides spectral efficiency predictions that closely match those obtained from tighter bounds of the Shannon ergodic rates, including the coherent decoding bound [1, 7].
Despite its broad success, the UatF bound has important limitations that restrict its general applicability. For example, it replaces the instantaneous effective channel by its mean and treats deviations as uncorrelated noise [3, Ch. 4.2], so it may lead to trivial bounds in scenarios where the beamformer remains fixed across multiple samples of a zero-mean channel. More specifically, under such condition, the numerator of the effective signal-to-interference-plus-noise ratio (SINR) expression in [3, Theorem 4.4] vanishes, which in turn yields a zero lower bound on the achievable rate. Furthermore, even if the UatF bound is nontrivial, it can severely underestimate the achievable rates when simple beamformers are employed, and its predictions can change even with a simple scaling of the samples of the beamforming vectors, an operation that should not affect the spectral efficiency if the decoder is aware of the scaling factor [3, Ch. 4.2.1]. For these reasons, there is a need to develop techniques that enable the derivation of power control algorithms based on alternative bounds. From an optimization standpoint, using known tight bounds is challenging because they often require computing the expectation of a logarithmic function involving fractions containing random variables (beamformers, channel estimates, etc.) in both the numerator and denominator [3, Theorem 5.1].
Against this background, we demonstrate that, in networks where the beamforming strategy is independent of the transmit power vector (the general case will be considered elsewhere), alternatives to the UatF bound – e.g., the coherent decoding bound and the upper bound in [8, Proposition 1] – can be employed in tractable optimization problems, provided we accept that expectations are approximated via Monte Carlo sampling, as also done in power control problems based on the UatF bound [2, 8, 7, 9]. Nevertheless, we avoid additional approximations, such as replacing the expectation of a ratio with the ratio of expectations.
To further contextualize our contribution, we note that many existing power control algorithms largely fall into two classes: (i) per-sample schemes, which solve an optimization problem for each channel sample [10, 11], but suffer from poor scalability in distributed MIMO systems; and (ii) statistical-level schemes, which solve a single optimization problem per channel distribution [1, 2], but rely on the UatF bound and thus inherit its limitations. Our framework builds upon the foundations of both classes to address these drawbacks. In particular, we use this framework to derive a fixed point algorithm – built on the iterations introduced to the wireless literature in [11] – that provably converges to globally optimal solutions of (weighted) max-min problems using an achievable rate bound that exploits instantaneous CSI at the decoder.
II Mathematical preliminaries
In this section, we clarify notation, review relevant results from the literature, and establish the mathematical framework that will serve as the foundation for deriving power control schemes in the next section.
II-A Notation and definitions
We use the convention that the set $\mathbb{N}$} of natural numbers does not include zero. The sets of nonnegative and positive reals are denoted by, respectively, ${\mathbb{R}}_{+}:=~[0,\infty[$ and ${\mathbb{R}}_{++}:=~]0,\infty[$. The elements of a vector ${\bm{x}}\in{\mathbb{R}}^{N}$ are denoted either as $(x_{1},\ldots,x_{N})$ or as $x_{1},\ldots,x_{N}]^{t}$, where $(\cdot)^{t}$ denotes the transpose operator. We adopt the latter notation whenever we need to emphasize that vectors should be interpreted as column vectors, ensuring that subsequent equations follow conventions commonly used in the wireless literature. Inequalities involving vectors are understood coordinate-wise. A norm $\|\cdot\|$ on ${\mathbb{R}}^{N}$ is monotone if $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{+}^{N})~{\bm{x}}\leq{\bm{y}}\Rightarrow\|{\bm{x}}\|\leq\|{\bm{y}}\|$. The standard $l_{1}$, $l_{2}$, and $l_{\infty}$ norms (all of which are monotone on ${\mathbb{R}}_{+}^{N}$) are denoted by $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, and $\|\cdot\|_{\infty}$, respectively. A sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ of vectors in ${\mathbb{R}}^{N}$ is said to converge to $\bm{x}$$^{\star}$ if $\lim_{n\to\infty}\|{\bm{x}}_{n}-{\bm{x}}^{\star}\|=0$ for some norm $\|\cdot\|$ (and, hence, for every norm on ${\mathbb{R}}^{N}$ because of the equivalence of norms in finite-dimensional spaces). To avoid potential misunderstandings, we say that $f:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{X}\subset{\mathbb{R}}^{N}$ and $\mathcal{Y}\subset{\mathbb{R}}^{M}$, is a function only if $M=1$ or a mapping if $M\geq 1$ (NOTE: for $M=1$, then $f$ is both a function and a mapping). The set of fixed points of a mapping $f:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{Y}\subset\mathcal{X}$, is denoted by $\mathrm{Fix}(f):=\{{\bm{x}}\in\mathcal{X}\mid f({\bm{x}})={\bm{x}}\}$. A mapping $f:{\mathbb{R}}^{N}\to{\mathbb{R}}^{M}$ is concave (respectively, convex) if each coordinate function is concave (respectively, convex).
The following technical result can be established with arguments analogous to those in [12, Lemmas 2 and 3][13], in combination with the standard result shown in [14, Theorem 10.3][15, Proposition 1].
Fact 1.
Let $(c_{1},\ldots,c_{N},\sigma)\in{\mathbb{R}}_{+}^{N+1}$ be a nonnegative vector such that $c_{1}>0$ and $\sigma>0$. Then the function
$f:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto\dfrac{x_{1}}{\log(1+s({\bm{x}}))},$
where $s:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto{c_{1}x_{1}}/{(\sum_{n=1}^{N}c_{n}~x_{n}+\sigma)}$, is concave. Furthermore, $f$ can be continuously extended to the domain ${\mathbb{R}}_{+}^{N}$, and this continuous extension, denoted by $\bar{f}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$, is a concave function satisfying $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~\bar{f}({\bm{x}})\geq\bar{f}({\bm{0}})=\sigma/c_{1}>0$.
A probability space is denoted by $(\Omega,\mathcal{F},\mathsf{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ is the event space, and $\mathsf{P}$ is the probability measure. The next result is a direct implication of [16, Proposition 2.16].
Fact 2.
Given a probability space $(\Omega,\mathcal{F},\mathsf{P})$ and a scalar $\delta\geq 0$, assume that a random variable $X:\Omega\to{\mathbb{R}}$ satisfies $X>\delta$ almost surely (a.s.) (or, equivalently, $X(\omega)>\delta$ for almost every $\omega\in\Omega$) and $E[X(\omega)]:=\int X(\omega)\mathsf{P}(\mathrm{d}\omega)$ is well-defined (i.e., $E[X(\omega)]\in{\mathbb{R}}_{+}$). [^1] We adopt the stochastic-programming notation $E[X(\omega)]$ rather than $E[X]$ to avoid ambiguity when deterministic decision variables are present. This distinction is especially important later in (5) and (8), where the notation makes explicit that the power vector is not a decision variable for the beamformers, which are random vectors. Then $E[X(\omega)]>\delta$.
We can now proceed with the contributions of this section.
II-B Framework for Power Control with Tight Ergodic Rate Bounds
The remainder of this section has two main objectives. First, we demonstrate that the results in [11] continue to hold for a slight variation of the concept of standard interference mappings originally introduced in the wireless literature in [10] (see Proposition 1). Second, we establish a connection between these results and a general class of functions that includes, as special cases, those commonly arising in weighted max-min optimization problems involving tight bounds on Shannon (ergodic) rates (Proposition 2). This latter result lays the groundwork for the subsequent sections, which explore further connections to concrete problems in the wireless domain. We begin by introducing the following variant of standard interference mappings:
Definition 1.
We say that $f:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}$ is a monotonic, scalable, and positive (MSP) function if the following properties hold:
  • (i)
    [monotonicity] $(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})(\forall{\bm{y}}\in{\mathbb{R}}_{++}^{N})~{\bm{x}}\leq{\bm{y}}\Rightarrow f({\bm{x}})\leq f({\bm{y}})$;
  • (ii)
    [scalability] $(\forall\alpha>1)(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~f(\alpha{\bm{x}})<\alpha f({\bm{x}})$; and
  • (iii)
    [positivity (bounded away from 0)] $\inf_{{\bm{x}}\in{\mathbb{R}}_{++}^{N}}f({\bm{x}})>0$.
Likewise, we say that a mapping $T:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(f_{1}({\bm{x}}),\ldots,f_{N}({\bm{x}}))$ is an MSP mapping if the coordinate functions $(f_{n}:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++})_{n\in\{1,\ldots,N\}}$ are MSP functions.
The key distinction between Definition 1 and the original standard interference functions in [10] lies in the domain of the mappings. Standard interference functions restricted to the positive cone ${\mathbb{R}}_{++}^{N}$ are MSP functions by definition, and both classes of functions are continuous on ${\mathbb{R}}_{++}^{N}$ [17]. However, the continuous extensions of MSP functions to the nonnegative cone ${\mathbb{R}}_{+}^{N}$, which are guaranteed to exist [17][18, Theorem 5.1.5], are not standard interference functions in general. Without additional assumptions, these existing results and the positivity property in Definition 1 guarantee only the preservation of positivity, monotonicity, and weak scalability on ${\mathbb{R}}_{+}^{N}\setminus{\mathbb{R}}_{++}^{N}$– i.e., the strict inequality in the scalability condition has to be relaxed to a weak inequality in general. For example, consider the continuous extension $\bar{f}:{\mathbb{R}}^{2}_{+}\to{\mathbb{R}}_{++}$ of the function $f:{\mathbb{R}}^{2}_{++}\to{\mathbb{R}}_{++}:(x_{1},x_{2})\mapsto 1+x_{1}+\sqrt{x_{2}}+\max\{x_{1}-1,0\}$ to the domain ${\mathbb{R}}_{+}^{2}$, which is just the function $f$ itself with its domain extended to ${\mathbb{R}}_{+}^{2}$. It satisfies all properties in Definition 1 on ${\mathbb{R}}_{++}^{2}$, but, for ${\bm{x}}=(1,0)\notin{\mathbb{R}}_{++}^{2}$, we do not have scalability because $(\forall\alpha>1)~\bar{f}(\alpha{\bm{x}})=\alpha\bar{f}({\bm{x}})$.
[…]
Although the above aspects might prevent the direct application of some results for standard interference functions to MSP functions – because, as discussed above, the set of continuous standard interference functions restricted to the positive cone is a proper subset of MSP functions [^2] In turn, MSP functions are a proper subclass of order-preserving subhomogeneous functions [18, 20]. –, many key properties remain analogous. In particular, the class of MSP functions is closed under addition, positive scaling, and finite maxima and minima, just as in the case of standard interference functions. Furthermore, any MSP mapping $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ is contractive in the complete metric space $({\mathbb{R}}_{++}^{N},d_{\mathrm{T}})$ [18, Lemma 2.1.7] (but not necessarily a Lipschitz contraction), where $d_{\mathrm{T}}:{\mathbb{R}}_{++}^{N}\times{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}$ is Thompson’s metric defined in [18, p. 30], so $T$ has at most one fixed point. If $\mathrm{Fix}(T)\neq\emptyset$, arguments similar to those described in [10] to prove convergence of the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via the fixed point iterations $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=T({\bm{x}}_{n})$, with ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, can be used for MSP mappings. In addition, the following property, which is crucial for the power control algorithms derived in the next section, can be established using the same arguments in [11], and they are therefore omitted.
Proposition 1.
Let $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ be an MSP mapping. Given a monotone norm $\|\cdot\|$, there exists exactly one tuple $(\gamma^{\star},{\bm{x}}^{\star})\in{\mathbb{R}}_{++}\times{\mathbb{R}}_{++}^{N}$ solving the so-called conditional (nonlinear) eigenvalue problem:
$\text{ Find }(\gamma,{\bm{x}})\in{\mathbb{R}}_{++}\times{\mathbb{R}}_{++}^{N}\text{ s.t. }T({\bm{x}})=\gamma{\bm{x}},~\|{\bm{x}}\|=1.$ (1)
Furthermore, the solution $(\gamma^{\star},{\bm{x}}^{\star})$ to (1) (called the conditional eigenpair) is such that the conditional eigenvalue $\gamma^{\star}$ satisfies $\gamma^{\star}=\|T({\bm{x}}^{\star})\|$ and the conditional eigenvector $\bm{x}$$^{\star}$ is the unique fixed point ${\bm{x}}^{\star}\in\mathrm{Fix}(G)$ of the mapping $G:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(1/\|T({\bm{x}})\|)T({\bm{x}})$. In addition, given ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=G({\bm{x}}_{n})$ converges to $\bm{x}$$^{\star}$.
The above result is especially relevant to the following class of functions:
Proposition 2.
Given a probability space $(\Omega,\mathcal{F},\mathsf{P})$, assume that $g:{\mathbb{R}}_{++}^{N}\times\Omega\to{\mathbb{R}}_{++}$ and $h:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}$ are functions such that, for almost every $\omega\in\Omega$, the function $q_{\omega}:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto z({\bm{x}},\omega)$, where
$z:{\mathbb{R}}_{++}^{N}\times\Omega\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto h({\bm{x}})/g({\bm{x}},\omega),$
is an MSP function. Further assume that there exists a random variable $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$, we have $z({\bm{x}},\omega)\geq\delta(\omega)>0$ for almost every $\omega\in\Omega$, and that the expectations $E[(\delta(\omega))^{-1}]$ and $E[g({\bm{x}},\omega)]:=\int g({\bm{x}},\omega)~\mathsf{P}(\mathrm{d}\omega)\geq 0$ are well defined for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ (i.e., $E[(\delta(\omega))^{-1}]\in{\mathbb{R}}_{++}$ and $(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~E[g({\bm{x}},\omega)]\in{\mathbb{R}}_{+}$). Then
$f:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto\dfrac{h({\bm{x}})}{E[g({\bm{x}},\omega)]}$ (2)
is an MSP function.
Proof.
(Monotonicity) Fix two vectors ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ and ${\bm{y}}\in{\mathbb{R}}_{++}^{N}$ such that ${\bm{x}}\leq{\bm{y}}$. For almost every $\omega\in\Omega$, $q_{\omega}$ is an MSP function by assumption, and thus we have the following chain of inequalities:
$\dfrac{1}{z({\bm{x}},\omega)}=\dfrac{g({\bm{x}},\omega)}{h({\bm{x}})}\geq\dfrac{1}{z({\bm{y}},\omega)}=\dfrac{g({\bm{y}},\omega)}{h({\bm{y}})}>0.$ (3)
Monotonicity of expectations and Fact 2 yield $1/f({\bm{x}})=E[g({\bm{x}},\omega)]/h({\bm{x}})\geq 1/f({\bm{y}})=E[g({\bm{y}},\omega)]/h({\bm{y}})>0$, which implies monotonicity of $f$.
(Scalability) Fix $\alpha>1$ and ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ arbitrarily. Using scalability and positivity of $q_{\omega}$ for almost every $\omega\in\Omega$, we deduce
$\dfrac{1}{z(\alpha{\bm{x}},\omega)}=\dfrac{g(\alpha{\bm{x}},\omega)}{h(\alpha{\bm{x}})}>\dfrac{1}{\alpha~z({\bm{x}},\omega)}=\dfrac{g({\bm{x}},\omega)}{\alpha h({\bm{x}})}>0.$ (4)
Taking expectations on both sides of the inequality, and using Fact 2 by setting $X:\Omega\to{\mathbb{R}}:\omega\mapsto 1/z(\alpha{\bm{x}},\omega)-1/(\alpha~z({\bm{x}},\omega))>0$, we obtain $1/f(\alpha{\bm{x}})=E[g(\alpha{\bm{x}},\omega)]/h(\alpha{\bm{x}})>E[g({\bm{x}},\omega)]/(\alpha h({\bm{x}}))=1/(\alpha~f({\bm{x}}))>0$, which implies scalability of $f$.
(Positivity) For almost every $\omega\in\Omega$, the function $q_{\omega}$ is an MSP function by assumption, so it can be continuously extended to the domain ${\mathbb{R}}_{+}^{N}$ while preserving monotonicity [18, Theorem 5.1.5]. We denote this continuous extension of $q_{\omega}$ to the domain ${\mathbb{R}}_{+}^{N}$ by $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$. Let $({\bm{y}}_{n})_{n\in{\mathbb{N}}}$ be a sequence in ${\mathbb{R}}_{++}^{N}$ converging to $\bm{0}$}. Hence, from the assumptions of the proposition, there exists $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for almost every $\omega\in\Omega$, we have
$(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~0<\delta(\omega)\leq\inf_{{\bm{y}}\in{\mathbb{R}}_{++}^{N}}{z}({\bm{y}},\omega)=\inf_{{\bm{y}}\in{\mathbb{R}}_{++}^{N}}\bar{q}_{\omega}({\bm{y}})\\ \leq\lim_{n\to\infty}\bar{q}_{\omega}({\bm{y}}_{n})=\bar{q}_{\omega}({\bm{0}})\stackrel{{\scriptstyle(a)}}{{\leq}}\bar{q}_{\omega}({\bm{x}})=z({\bm{x}},\omega)\in{\mathbb{R}}_{++},$
where (a) follows from monotonicity of $\bar{q}_{\omega}$, as already established above. Taking reciprocals, computing expectations, and recalling that $E[(\delta(\omega))^{-1}]<\infty$ by assumption, we verify that
$(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~\infty>E\left[\dfrac{1}{\delta(\omega)}\right]\geq E\left[\dfrac{1}{{z}({\bm{x}},\omega)}\right]=\dfrac{1}{f({\bm{x}})}>0,$
which implies the desired result $\inf_{{\bm{x}}\in{\mathbb{R}}_{++}^{N}}f({\bm{x}})\geq(E\left[{1}/{\delta(\omega)}\right])^{-1}>0$. ∎
III Power control in multi-user MIMO
We consider the uplink of a general multi-user MIMO network with $L\in{\mathbb{N}}$ access points, each equipped with $M\in{\mathbb{N}}$ antennas. Thus, the total number of antennas on the access point side is $K:=LM$. The set of access points is denoted by $\mathcal{A}:=\{1,\ldots,L\}$. The system also includes $N\in{\mathbb{N}}$ single-antenna users, which are indexed by the set $\mathcal{U}:=\{1,\ldots,N\}$. The channel and beamforming vectors for user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ are random (column) vectors denoted by, respectively, ${\bm{h}}_{u,a}:\Omega\to\mathbb{C}^{M}$ and ${\bm{v}}_{u,a}:\Omega\to\mathbb{C}^{M}$, where $(\Omega,\mathcal{F},\mathsf{P})$ is the probability space. The transmit power of user $u\in\mathcal{U}$ is denoted by $p_{u}\in{\mathbb{R}}_{++}$. Without any loss of generality, we assume that $(\forall\omega\in\Omega)~{\bm{v}}_{u,a}(\omega)={\bm{0}}$ if access point $a\in\mathcal{A}$ does not participate in the processing of the signal transmitted by user $u\in\mathcal{U}$, and that $\|{\bm{v}}_{u}(\omega)\|_{2}=1$ for almost every $\omega\in\Omega$ and every $u\in\mathcal{U}$. The aggregated channel vector and aggregated beamforming vector for user $u\in\mathcal{U}$ are denoted by, respectively, ${\bm{h}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{h}}_{u,1}^{t}(\omega),\ldots,{\bm{h}}^{t}_{u,L}(\omega)]^{t}$ and ${\bm{v}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{v}}^{t}_{u,1}(\omega),\ldots,{\bm{v}}^{t}_{u,L}(\omega)]^{t}$.
The above model covers many common technologies in the literature, including multi-cell massive MIMO networks as defined in [3] and user-centric cell-less networks with different levels of cooperation as defined in [2]. For brevity and notational simplicity (see Remark 3 below), we evaluate the utility (performance) of user $u\in\mathcal{U}$ as a function of the power allocation ${\bm{p}}=[p_{1},\ldots,p_{N}]^{t}\in{\mathbb{R}}_{++}^{N}$ via the following well-known achievable rate under perfect channel state information (CSI) [8, Proposition 1][21, Lemma 1]:
$r_{u}:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{+}:{\bm{p}}\mapsto E\left[\log(1+s_{u}({\bm{p}},\omega))\right],$ (5)
where
$\begin{array}[]{rcl}s_{u}:{\mathbb{R}}_{++}^{N}\times\Omega&\to&{\mathbb{R}}_{+}\\ ({\bm{p}},\omega)&\mapsto&\dfrac{p_{u}|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}}{\sum_{k\in\mathcal{U}\setminus\{u\}}p_{k}|{\bm{h}}_{k}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}+\sigma^{2}},\end{array}$ (8)
can be interpreted as the instantaneous SINR function and $\sigma^{2}\in{\mathbb{R}}_{++}$ is the noise power. Important aspects of the model we consider are discussed below.
Remark 1.
The beamformers in (5) are random vectors, so they are allowed to change for each channel realization. However, as clear from the notation, they should not depend on the power vector $\bm{p}$}, the argument of the functions $(r_{u})_{u\in\mathcal{U}}$. Prominent beamforming strategies satisfying this assumption include conventional conjugate beamforming, (regularized) zero-forcing beamforming, and MMSE-like beamformers such as those simulated in [9] and the references therein. The general case will be considered elsewhere.
Remark 2.
The algorithms derived later require the computation of the expectation in (5). In practice, this expectation can be approximated via Monte Carlo sampling, which requires statistical knowledge of the system. Although this assumption may appear stringent, statistical information, whether complete or partial, and Monte Carlo sampling are likewise required in power control schemes based on the UatF bound whenever existing channel models are involved [1, 2, 3] and additional approximations are avoided. In many applications, computing expectations via empirical sample means is possible; e.g., in applications such as network planning and in the creation of datasets for training neural networks for power control [22]. Furthermore, in online applications with iterative power updates during data transmission, SINR samples can be obtained directly from pilot signals; e.g., from demodulation reference signals (DMRS) in 5G. A rigorous analysis of how finite-sample noise affects convergence is deferred to future studies. However, Sect. IV provides empirical evidence that its impact on the performance of the proposed algorithms can be limited, mirroring empirical results widely reported for UatF-based power allocation schemes requiring Monte Carlo sampling.
Remark 3.
The utility in (5) is often called the “genie-aided” or “optimistic” upper bound for the achievable ergodic rates (when interference is treated as noise) because it assumes perfect CSI at the receiver [2, p. 328]. Nevertheless, the results we derive can be readily extended to other bounds, such as the coherent decoding lower bound in [2, Theorem 5.1][3, Theorem 4.1]. While this bound uses channel estimates – instead of perfect channels – and introduces an “error covariance” term in the denominator of the expectation in (5), the subsequent analysis in this study applies to it in a line-by-line manner with only straightforward changes.
We now establish basic properties of the utility functions $(r_{u})_{u\in\mathcal{U}}$ that will allow us to derive power control algorithms based on the fixed point iterations described in the previous section.
For every user $u\in\mathcal{U}$, we assume that $r_{u}({\bm{p}})>0$ if ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$ $r_{u}$ is defined in (5)]. Otherwise, the user can be excluded from consideration because such a user cannot be served by the network. Therefore, denoting by $\gamma_{u}\in{\mathbb{R}}_{+}$ the utility of user $u\in\mathcal{U}$ for a given power ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$, we deduce
$0<\gamma_{u}=r_{u}({\bm{p}})\Leftrightarrow p_{u}={\gamma_{u}}f_{u}({\bm{p}}),$
where
$(\forall u\in\mathcal{U})~f_{u}:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}:{\bm{p}}\mapsto\dfrac{p_{u}}{r_{u}({\bm{p}})}.$ (9)
The next result establishes a fundamental connection among the utility in (5), the function in (9), and Proposition 2:
Proposition 3.
For every user $u\in\mathcal{U}$ and power vector ${\bm{p}}\in{\mathbb{R}}^{N}_{++}$, assume that the expectation in (5) is well defined, $E[|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}]<\infty$, and $|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|\neq 0$ for almost every $\omega\in\Omega$. Then $f_{u}:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}$ in (9) is an MSP function.
Proof.
Fix $u\in\mathcal{U}$ arbitrarily. Fact 1 and the assumption $|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|\neq 0$ (a.s.) show that, for almost every $\omega\in\Omega$, the function
$q_{\omega}:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}:{\bm{p}}\mapsto p_{u}/\log(1+s_{u}({\bm{p}},\omega)),$
is concave $s_{u}$ is defined in (8)], and it has a continuous concave extension $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}$ satisfying $(\forall{\bm{p}}\in{\mathbb{R}}_{++}^{N})~q_{\omega}({\bm{p}})\geq\bar{q}_{\omega}({\bm{0}})=\sigma^{2}/|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}>0$. Now apply [15, Proposition 1] to $\bar{q}_{\omega}$ to conclude that ${q}_{\omega}$, the restriction of $\bar{q}_{\omega}$ to the domain ${\mathbb{R}}_{++}^{N}$, is an MSP function. The desired result now follows as an immediate application of Proposition 2. ∎
We now have all the necessary background to pose and solve a power control problem that has the objective of maximizing the minimum (weighted) achievable rate while satisfying the power constraint $\|{\bm{p}}\|\leq p_{\mathrm{max}}$, where $\|\cdot\|$ is a given monotone norm (typically the $l_{\infty}$-norm in the uplink) and $p_{\mathrm{max}}>0$ the maximum transmit power. Formally, we consider the following optimization problem:
$\begin{array}[]{rl}\text{maximize}_{{\bm{p}}\in{\mathbb{R}}_{++}^{N}}&\min_{u\in\mathcal{U}}\alpha_{u}^{-1}~r_{u}({\bm{p}})\\ \text{s.t.}&\|{\bm{p}}\|\leq p_{\textrm{max}},\end{array}$ (12)
where $(\alpha_{1},\dots,\alpha_{N})\in{\mathbb{R}}_{++}^{N}$ corresponds to the weights (priorities) assigned to users, and $r_{u}$ is the achievable rate of user $u\in\mathcal{U}$ in (5). In light of (9), the epigraph form of the above problem is given by:
$\begin{array}[]{rl}\text{maximize}_{(\gamma,{\bm{p}})\in{\mathbb{R}}_{++}\times{\mathbb{R}}_{++}^{N}}&\gamma\\ \text{s.t.}&{\bm{p}}\geq\gamma T({\bm{p}})\\ &\|{\bm{p}}\|\leq p_{\textrm{max}},\end{array}$ (16)
where $T:{\mathbb{R}}_{++}^{N}\mapsto{\mathbb{R}}_{++}^{N}:{\bm{p}}\mapsto[\alpha_{1}f_{1}({\bm{p}}),\ldots,\alpha_{N}f_{N}({\bm{p}})]^{t}$, with $(f_{u})_{u\in\mathcal{U}}$ as defined in (9), is an MSP mapping as an immediate consequence of Proposition 3.
We can use the same arguments in [5, Proposition 2] with standard interference functions replaced with MSP functions to show that, although the solution to (16) (and, hence, (12)) may not be unique, there always exists a solution achieving equality in all constraints; i.e., there exists a solution $(\gamma^{\star},{\bm{p}}^{\star})$ satisfying $(\gamma^{\star})^{-1}{\bm{p}}^{\star}=T({\bm{p}}^{\star})$ and $(1/p_{\text{max}})\|{\bm{p}}^{\star}\|=1$. These equalities establish that $(\gamma^{\star})^{-1}$ and $\bm{p}$$^{\star}$ are, respectively, the conditional eigenvalue and conditional eigenvector of the mapping $T$ and the monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|_{\star}:=(1/p_{\text{max}})\|{\bm{p}}\|$. Therefore, as a corollary of Proposition 1, Proposition 3, and [5, Proposition 2] (after adapting the proof of this last proposition by substituting MSP functions with standard interference functions), we obtain the following result, which shows a simple fixed point method that provably converges to the solution to (12) with minimum total power:
Corollary 1.
Let the assumptions in Proposition 3 be valid. Denote by $\mathcal{S}\neq\emptyset$ the set of solutions to (12) and let $\|\cdot\|$ be the monotone norm in (12). Given ${\bm{p}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence generated $({\bm{p}}_{n})_{n\in{\mathbb{N}}}$ via $(\forall n\in{\mathbb{N}})~{\bm{p}}_{n+1}=(p_{\mathrm{max}}/\|T({\bm{p}}_{n})\|)T({\bm{p}}_{n})$ converges to the power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$ that solves the optimization problem
$\text{minimize}_{{\bm{p}}\in\mathcal{S}}\|{\bm{p}}\|_{1}.$
IV Simulations
We consider a cell-less network with $L=16$ access points, each equipped with a four-element uniform linear array with half-wavelength spacing ($\lambda/2$), where $\lambda=c/f_{\mathrm{c}}$ is the wavelength, $f_{\mathrm{c}}$ is the carrier frequency, and $c$ is the propagation speed. The access points are placed uniformly over a $1000\times 1000~\mathrm{m}^{2}$ area. We distribute $N=25$ single-antenna users uniformly at random and apply a conventional wrap-around technique to emulate an unbounded service area. The user-access point antenna height difference is $\Delta h=11$ m. The large-scale fading (path loss) $\beta_{u,a}$ between user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ follows the COST-231 Walfish–Ikegami UMi model [23, Sect. 5.2]:
$\beta_{u,a}=-35.4+20\log_{10}(f_{\mathrm{c}})+26\log_{10}\left(\frac{d_{u,a}}{1~\text{m}}\right)+F_{u,a}~[\text{dB}],$
where $d_{u,a}$ is the 3D distance and $F_{u,a}\sim\mathcal{N}(0,\sigma_{\mathrm{sf}}^{2})$ models shadow fading (in dB). Small-scale fading follows a spatially correlated Rician model. All beamformers operate on channel estimates – rather than the true channels – obtained using the estimation procedure described in [7]. For pilot assignment and formation of user-centric cooperation clusters in the cell-less network, we use the algorithm described in [2, Algorithm 4.1]. The main simulation parameters are listed in Table I, and we refer the readers to [7] for further details on the simulation.
Parameter Value
Network area 10001000 × 10001000 $\text{m}^{2}$
Number of APs $L=16$
Number of users $N=25$
Number of antennas per access point $M=4$
Bandwidth $B=20$ MHz
Carrier frequency $f_{c}=3.7$ GHz
Maximum uplink transmit power $p_{\text{max}}=200$ mW
($\|\cdot\|_{\infty}$ used as the
monotone norm in (12))
Coherence block symbols $\tau_{\mathrm{c}}=200$
Number of pilot symbols $\tau_{\mathrm{p}}=10$
Access point-user height difference $\Delta h=11$ m
Shadow fading (Line-of-Sight) $\sigma_{\text{sf}}=8$ dB
Antenna spacing $d=\lambda/2$
TABLE I: Simulation parameters
The performance evaluation considers uniform weights in (12) (i.e., $(\forall u\in\mathcal{U})~\alpha_{u}=1$) and centralized cell-less MMSE beamforming design with fixed (unitary) power weighting coefficients in the channel inversion stage [9, Sect. 3-D]. We estimate the ergodic rates by averaging across 500 Monte Carlo trials, each drawing an independent channel realization for a single, fixed network layout. In the proposed fixed point iterations, shown in Corollary 1, we use (5) as the utility function, which is called “optimistic ergodic rate” (OER) in the following (see Remark 3). As a baseline, we compute the OER bound with the solution to the standard max-min problem under the UatF bound, which admits a pseudo-closed-form expression [9, Prop. 4]. This approach mirrors established practice in the literature: if simple fixed point algorithms capable of solving (12) were unknown (as was the case prior to this study, to the best of our knowledge), we would consider the surrogate max-min power-control problem that uses the UatF bound instead of the target bound in (5), with the expectation that the surrogate solution would approximate the solution to (12).
Figure 1: Convergence of the worst-user rate (OER bound) under centralized MMSE beamforming. For reference, the minimum OER bound obtained with the solution to the UatF max-min problem is included.
Fig. 1 shows the convergence of the objective value for the problem in (12) using the fixed point algorithm in Corollary 1, and, for comparison, the objective value computed with the solution to the UatF max-min problem. We verify that the proposed fixed point algorithm shows negligible drift from finite-sample Monte Carlo estimation of expectations, indicating robustness to sampling noise. Furthermore, replacing the OER bound with the UatF surrogate in the max-min problem leads to a noticeable performance loss with respect to the OER bound. This result is hardly surprising: the proposed iterative algorithm provably converges to a global solution to (12), which uses the OER bound, the figure of merit considered in the simulations.
V Conclusion
The UatF bound has been widely used in network optimization because dealing with alternative information-theoretic (ergodic) rate bounds were viewed as difficult. However, we have shown a framework able to target these alternative bounds directly, giving rise to simple fixed point algorithms that provably converge to global optima of max-min power control problems in cellular and cell-less massive MIMO systems, among others. Simulations show that these algorithms deliver gains in regimes where the UatF bound is overly conservative. Finally, we remark that the possible gains can be extreme in scenarios where statistical (nonadaptive) beamforming is used under zero-mean channels, in which case the UatF is inapplicable because it provides trivial (zero) rate bounds.